OpenStudy (anonymous):

What is the magnitude and direction of if G(2, -2) and H(7, 6)?

OpenStudy (lgbasallote):

hint: \[\Large \text{magnitude} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\Large \text{direction} = \tan^{-1} \left(\frac{y_2 - y_1}{x_2 - x_1} \right)\] does that help?

OpenStudy (anonymous):

magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (anonymous):

kind of, does that mean the answer is the 3rd one

OpenStudy (lgbasallote):

i'll check

OpenStudy (lgbasallote):

yup right

OpenStudy (anonymous):

thanks :D

OpenStudy (lgbasallote):

welcome ^_^